OpenStudy (anonymous):

please solve -1 +2x - 6 -3x - 100 = -5x +1

OpenStudy (anonymous):

combine like terms. (-1-6-100)+(2x-3x)=-5x+1 -107-x=-5x+1 -x+5x=1+107 4x=108 x=108/4

OpenStudy (anonymous):

Thanks a lot